OpenStudy (anonymous):

(Complex Variable) Determine for which values of a and b you can assure that the following integrals converge:

OpenStudy (anonymous):

a) \[\int\limits_{-1}^{1} \frac{\cos(it)}{t^a}\] I have b) and c) but I want help with one and I'll see if I can solve the other ones.

OpenStudy (anonymous):

I think it converges for all a with Re(a)>-1

OpenStudy (anonymous):

Sorry, Re(a)<1, I didn't think it as a fraction.

OpenStudy (john_es):

It's an interesting problem, I have some doubts about it, but may be it should be, \[Re(a)\leq1\]?

OpenStudy (anonymous):

Does it converge when Re(a)=1?

OpenStudy (john_es):

I cannot assure you, but I would say yes (I think there should be a hypergeometric function around there)

OpenStudy (anonymous):

I also was doing the following integrals: \[\int\limits_{1}^{\infty} \frac{t^a}{i+t^b} dt\] And \[\int\limits_{0}^{1} \frac{t^a}{i+t^b}\] The second one I think it's integrable if the denominator never tends to zero (I think it could be integrable even if it went to zero but it would require more analysis). So b real part can never be 1 if sin(ln(t)Im(b))=-1 In the other one I think that Re(a)<Re(b)+1 otherwise the integral I'm using to delimit it will not converge.